Say that, given the wonderful weather in the city, many workers from other states move to...

  1. Say that, given the wonderful weather in the city, many workers from other states move to Los Angeles to look for work. As a result

    Real wages W/P will remain unchanged

    Real wages W/P will increase in Los Angeles

    Real wages W/P will decrease in Los Angeles

    It is not possible to say whether real wages will increase or decrease, it depends on how many workers move to Los Angeles

    None of the above/below

QUESTION

  1. Say that firms pay a wage that is above the equilibrium wage to induce workers to exert more effort. The resulting unemployment is

    inexistent, there can’t be any unemployment when firms pay higher wages

    the natural rate of unemployment

    structural unemployment

    frictional unemployment

    None of the above/below

QUESTION

  1. Which of the following is not part of US GNP?

    a.

    The wages earned by a US national from working at Coffee Bean

    b.

    The net rent earned by a US national from renting one of his properties in Canada

    c.

    The net profits made by a US national from his start-up company located in Riverside

    d.

    None of the above/below

    e.

    The wages paid by the University of California to one of his professors, who is a national of the United Kingdom

QUESTION

  1. Say that the government announces that it will increases taxes on deposits. In response to this policy, households in the economy decide to withdraw some of their deposits from their banks and hold cash instead. Which of the following is true:

    None of the above/below is true

    The money supply will decrease

    There is no effect on the money supply, only the Central Bank can change the money supply

    Whether the money supply increases or decreases overall depends on the size of the tax

    The money supply will increase

QUESTION

  1. Suppose that Japan’s Central Bank announces it wants to target an inflation rate of at least 2%, and that real GDP growth in the economy is 1%. Assume that velocity is constant. Then money growth must be at least

    2%

    4%

    1%

    None of the above/below

    3%

QUESTION

  1. Katherine consumes only cookies and chocolates. In year 1, cookies cost $3 each, chocolates cost $4 each, and Katherine buys 10 cookies and no chocolates. In year 2, cookies cost $4 each, chocolates cost $3 each, and Katherine buys 10 chocolates and no cookies. Taking Katherine’s consumption as the typical consumption basket, the CPI index in year 2 (using year 1 as the base year) is:

    75

    66.67

    100

    133.33

    None of the above/below
Answer #1

1. The immigration of workers in a state reduces the real wage since the supply of labour increase.

Answer: Real wage W/P will decrease in Loss Angeles.

2. At a given wage rate if the quantity of labour supplied exceeds the quantity of labour demanded, this unemployment is known as structural unemployment. There is a mismatch between number of people who want to work and the number of jobs available.

Answer: Structural unemployment.

3. GDP is the money value of goods and services produced within the national boundary of a country. But GNP includes the money value of goods and services produced domestically plus income earned by the residents of the country from the rest of the world minus the income earned inside the country by foreign residents. Here the income earned by the professor of United Kingdom to be deducted in the calculation of GNP

Answer: e. The wages paid by the University of California to one of its professors who is a national of the United Kingdom.

4. When the people withdraw money from their accounts, the bank reserve decreases. This will reduce the issue of loans from the banks and the money supply in the economy.

Answer: The money supply will decrease.

5. According to the quantity equation MV=PT. MV is the supply of money and PT is the demand for money. Total money supply is M×V and total demand for money is P×T. If V is constant the total money supply M=P×T.

If inflation is 2% and GDP (T) 1% then the money growth will be 2×1=2.

Answer: 2%

6. Consumer Price Index (CPI)= Current year cost of basket/Base year cost of basket ×100

Cost of basket (Year 1)

Cookies- $3×10= $30

Chocolates -$4×10= $40

Cost of basket in year 1= $70

Cost of basket (Year 2)

Cookies- $4×10= $40

Chocolates -$3×10= $30

Cost of basket in year 7= $70

CPI= 70/70×100=100

Answer: 100.